(2 points) Solve the initial value EN dy - 2xy = 6x, yO) = 4
Answer #1

dy -247 = 6x , yco= 4 dy - 6x + 2xy = 22 (3+y) [C=7] 1973 dy = (2xdX en 19+3) = x² + ene But y(0)=4 enz= otenc = en 19+3) = x